Twice the larger of two numbers is three more than five times the smaller and the sum of four times the larger and three times the smaller is 71. What are the numbers?
Answer by mananth(16949)

Twice the larger of two numbers is three more than five times the smaller and the sum of four times the larger and three times the smaller is 71. What are the numbers?

let x >y
2x=5y+3
2x-5y=3.................(1)

4x+3y=71................(2)
2 x -5 y = 3 .............1

4 x + 3 y = 71 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 5
6 x -15 y = 9
20 x + 15 y = 355
Add the two equations
26 x = 364
/ 26
x = 14
plug value of x in (1)
2 x -5 y = 3
28 -5 y = 3
-5 y = 3 -28
-5 y = -25
y = 5
14 &5